# DriveGen3D: Boosting Feed-Forward Driving Scene Generation with Efficient Video Diffusion

### v1

@@ Abstract
We present DriveGen3D, a novel framework for generating high-quality and highly controllable dynamic 3D driving scenes that addresses critical limitations in existing methodologies. Current approaches to driving scene synthesis either suffer from prohibitive computational demands for extended temporal generation, focus exclusively on prolonged video synthesis without 3D representation, or restrict themselves to static single-scene reconstruction. Our work bridges this methodological gap by integrating accelerated long-term video generation with large-scale dynamic scene reconstruction through multimodal conditional control. DriveGen3D introduces a unified pipeline consisting of two specialized components: FastDrive-DiT, an efficient video diffusion transformer for high-resolution, temporally coherent video synthesis under text and Bird’s-Eye-View (BEV) layout guidance; and FastRecon3D, a feed-forward reconstruction module that rapidly builds 3D Gaussian representations across time, ensuring spatial-temporal consistency. Together, these components enable real-time generation of extended driving videos (up to $424\times 800$ at 12 FPS) and corresponding dynamic 3D scenes, achieving SSIM of 0.811 and PSNR of 22.84 on novel view synthesis, all while maintaining parameter efficiency.
Figure 1: DriveGen3D visualization. Comparison of GT videos and generated video for reconstruction.
1 Introduction
@@
The synthesis of 3D dynamic driving environments has emerged as a key research frontier in autonomous systems, driven by its wide-ranging applications in simulation, perception, and planning. While recent advances in video generation [1, 2, 3, 4, 5, 6, 7] and 3D scene reconstruction [8, 9, 10, 11, 12, 13, 14, 15, 16] have made substantial progress, a critical gap remains: the lack of an integrated and efficient framework that unifies long-horizon video synthesis and large-scale 3D scene reconstruction under multimodal control.
@@
Existing methodologies typically address either temporal coherence in video generation or spatial fidelity in scene reconstruction—but not both—often requiring high computational cost or suffering from limited scalability. For example, state-of-the-art diffusion-based models like MagicDriveDiT [2] can produce high-resolution driving sequences but require up to 30 minutes to generate a single $1600\times 900$ video, making them impractical for real-time use.
@@
From the reconstruction perspective, optimization-based approaches [17, 8, 9, 10] are similarly time-consuming, often needing 30 minutes per scene. While recent feed-forward methods [18, 19, 20, 11, 12, 13] have reduced reconstruction time to seconds, they remain limited in scale and rarely integrate with dynamic video generation pipelines.
@@
To bridge this gap, we propose DriveGen3D, an efficient and unified framework that integrates two specialized modules: FastDrive-DiT, an accelerated video diffusion transformer for high-resolution driving video generation, and FastRecon3D, a feed-forward reconstruction pipeline that builds dynamic 3D scenes from multi-view video frames in real time. FastDrive-DiT employs both diffusion step caching and quantized attention to reduce inference time by over $2\times$, while FastRecon3D leverages temporal-aware Gaussian splatting to produce high-fidelity reconstructions with minimal latency. Together, these components enable high-quality video generation and complete 3D reconstruction within 6 minutes, significantly outperforming prior methods in both efficiency and scalability, as shown in Figure 2.
Our key contributions are summarized as follows:
@@
  • We conduct extensive experiments demonstrating that DriveGen3D outperforms state-of-the-art baselines in both video quality and 3D reconstruction accuracy. Specifically, our framework achieves a reconstruction SSIM of 0.811 on generated inputs, surpassing the previous best (0.767) while maintaining competitive PSNR. Moreover, DriveGen3D reduces the end-to-end generation time to under 6 minutes, including video synthesis and 3D scene reconstruction, marking a substantial improvement in efficiency.
Figure 2: Performance vs. Efficiency. DriveGen3D achieves the highest SSIM (0.811) while significantly reducing generation time to 6 minutes—an 80% improvement over optimization-based and diffusion-based baselines—demonstrating superior video quality and real-time capability.
2 Related Work
2.1 Video generation for driving scene
@@
Recent advancements in street view generation and autonomous driving scene synthesis have significantly improved the fidelity and controllability of synthetic data. MagicDrive [1] introduces a framework for street view generation with diverse 3D geometry controls, such as camera poses and 3D bounding boxes, enhancing tasks like BEV segmentation and 3D object detection through cross-view consistency. MagicDriveDiT [2] extends this by addressing high-resolution, long video generation for autonomous driving, leveraging flow matching and spatial-temporal conditional encoding to achieve superior scalability and control. InfiniCube [21] focuses on unbounded dynamic 3D driving scene generation, combining scalable 3D representations with video models to ensure geometric and appearance consistency across large-scale scenes. DrivingDiffusion [3] synthesizes multi-view driving videos with precise layout control, ensuring cross-view and cross-frame consistency through a spatial-temporal diffusion framework. UniScene [22] unifies the generation of semantic occupancy, video, and LiDAR data, employing a progressive generation process to reduce complexity and improve downstream task performance. DriveDreamer [4] pioneers real-world-driven world models, using diffusion models to capture complex driving environments and enhance driving video generation and action prediction. Panacea [23] integrates a novel 3D attention and a two-stage generation pipeline to maintain coherence, supplemented by the ControlNet framework for meticulous control by the Bird’s-Eye-View (BEV) layouts. DriveDreamer-2 [5] integrates LLMs to generate user-defined driving videos, improving temporal and spatial coherence while surpassing state-of-the-art methods in video quality metrics like FID and FVD. Together, these works advance the field of autonomous driving by providing scalable, controllable, and high-fidelity synthetic data generation frameworks.
2.2 Reconstruction for driving scene
. The reconstruction of dynamic driving scenes [8, 17, 24, 25, 3, 26, 27, 28, 29, 30] has emerged as a critical task in autonomous systems and immersive environment modeling. Contemporary approaches predominantly leverage Gaussian splatting-based representations due to their inherent balance between rendering efficiency and geometric expressiveness. Early methodologies in this domain adopted optimization-based paradigms, exemplified by works such as StreetGaussian [8], DrivingGaussian [10], and HUGS [9]. These frameworks optimize scene representations per instance for specific street segments (typically under 100 meters in scale) through iterative refinement processes spanning approximately 30 minutes.
@@ 2.2 Reconstruction for driving scene
Recent advancements have shifted toward feed-forward architectures [18, 19, 20, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41] to enable rapid 3D reconstruction. Methods like PixelSplat [18], MVSplat [19], and DepthSplat [20] employ large pretrained networks to directly infer Gaussian parameters from multi-view inputs, reducing reconstruction time from minutes to seconds. Though these approaches demonstrate generalizability across scenes, they often sacrifice reconstruction fidelity in geometrically complex regions or under sparse observational constraints. Parallel innovations address the temporal dimension of driving scenes: InfiniCube [21] extends the Scube [42] framework to 3D street generation, disentangling dynamic vehicles from static backgrounds via hybrid voxel-video control mechanisms. Drive3R [13] adapts the Spann3R [31] architecture for per-frame 3D scene reconstruction through temporal consistency priors. DrivingRecon [12] mimics StreetGaussians’ pipeline but replaces iterative optimization with feed-forward prediction, achieving real-time capability at moderate resolutions. DrivingForward [11] enhances sparse-view reconstruction robustness by jointly learning pose estimation and depth prediction modules within its network architecture. These advancements collectively highlight two persistent limitations:
@@ 3.1 Overview
DriveGen3D, as illustrated in Figure 3, is an integrated 3D driving scene generation system composed of two key components: FastDrive-DiT for efficient long video generation, and FastRecon3D for feed-forward 3D scene reconstruction. The pipeline begins with FastDrive-DiT, which synthesizes high-resolution, temporally coherent driving videos under conditional guidance. These generated videos are then passed to FastRecon3D, which reconstructs dynamic 3D scenes in a feed-forward manner using temporal-aware Gaussian splatting. Together, these modules enable rapid and scalable 3D scene generation suitable for real-time simulation and autonomous driving applications.
The architecture of our efficient video diffusion module, dubbed FastDrive-DiT, integrates two targeted acceleration strategies to address the high inference cost of driving video generation. Diffusion steps acceleration: Given the slow inference speed of the driving video generation diffusion process, we employ TeaCache [43] to expedite the generation. In contrast to the original approach, we utilize TeaCache only for the conditional branch of the video generation model. This modification reduces the calculation cost of TeaCache by half, leading to a speedup of more than $2\times$. Quantized DiT: Considering the high cost of the attention part in the video generation model, we use SageAttention [44, 45] for the quantization of $Q,K,V$ during the inference procedure. We conduct inference profiling and apply SageAttention to all the attention blocks, we are able to save an additional 30 seconds of inference time.
The architecture design of FastRecon3D is similar to DrivingForward [11]. The generated image is first processed through an image encoder, after which it is fed into decoders to obtain camera pose estimation and depth information respectively. These derived parameters are subsequently combined with the encoded image features and fed into the Gaussian prediction network, which predicts per-frame Gaussian parameters for dynamic scene representation.
3.2 FastDrive-DiT
@@
Generating 3D driving scenes, especially in the autonomous driving domain, is notoriously time-consuming due to the multi-view nature of the data. The video generation step is particularly costly because of the underlying diffusion process. For instance, MagicDriveDiT can take up to 30 minutes to produce a video of resolution $1600\times 848\times 6\times 233$. To address this inefficiency, we propose FastDrive-DiT, an enhanced and lightweight video diffusion model built on MagicDriveDiT with targeted acceleration strategies.
@@
Diffusion steps acceraleration. To accelerate the generation of long and high-resolution videos in MagicDriveDiT, we integrate TeaCache[43], a training-free caching approach that enhances the inference speed of diffusion models. TeaCache leverages timestep embeddings to modulate noisy inputs, approximating the fluctuating differences in model outputs across timesteps. By introducing a rescaling strategy, it refines these differences, enabling efficient output caching with minimal computational overhead. Experiments show that TeaCache achieves up to 4.41× acceleration over Open-Sora-Plan while maintaining visual quality, making it an effective solution for balancing speed and performance in video generation.
@@
In the original TeaCache implementation, coefficients are computed from both the conditional and unconditional branches of the diffusion model. However, in our adaptation for MagicDriveDiT, we calculate coefficients exclusively from the conditional branch. This optimization not only reduces the forwarding time of TeaCache but also ensures no obvious degradation in performance, further enhancing the efficiency of the video generation process.
@@
3.3 FastRecon3D
While the aforementioned methods enable realistic driving scene generation, applications like simulation require complete 3D scene models. To enable rapid novel scene synthesis, we introduce FastRecon3D, a feed-forward reconstruction module that avoids costly optimization while preserving quality. Built upon temporal-aware Gaussian Splatting, FastRecon3D reconstructs each frame’s 3D geometry by leveraging both past and future contexts, enabling fast and consistent 3D scene generation.
@@ 3.3 FastRecon3D
$G_{i}^{t}=\{\mathcal{G}_{i}^{t}\}_{i=1}^{N_{t}}=\{(\mathbf{\mu}_{i},\mathbf{\Sigma}_{i},\mathbf{\alpha}_{i},\mathbf{c}_{i})\}_{i=1}^{N_{t}}$ (1)
Recursively reconstruction from videos. We modified DrivingForward [11], a state-of-the-art 3D reconstruction model for autonomous driving, to build the scene.
After the frame generation of the $t+1$ timestep, we extract all frames from time steps t1t-1 to $t+1$ to reconstruct the scene at time $t$. Formally, for each timestep $t$, given multi-view images $\{I_{i}^{t}\}_{i=1}^{N_{t}}$ and temporal neighbors $\{I_{i}^{t\pm\Delta}\}_{i=1}^{N_{t}}$, our model predicts Gaussian parameters $\mathcal{G}^{t}=\{\mu^{t},\Sigma^{t},\alpha^{t},c^{t}\}$ and save it as a 3D model:
$\mathcal{G}^{t}=\mathcal{F}\Bigl(\{I_{i}^{t-\Delta},\,I_{i}^{t},\,I_{i}^{t+\Delta}\}_{i=1}^{N_{t}}\Bigr).$ (2)
By leveraging both past and future context, this recursive reconstruction method effectively captures dynamic scene elements while maintaining high spatial fidelity. As a result, our approach produces complete 3D models in a matter of seconds, meeting the demanding requirements of simulation and other real-time applications without compromising on quality. The overall pipeline of DriveGen3D is outlined in Algorithm 1.
Algorithm 1 Overall pipeline of DriveGen3D.
1:Text prompt $\mathcal{T}$, BEV layout $\mathcal{B}$
2:3D driving scene $\mathcal{G}^{1:T}$
3:Stage 1: FastDrive-DiT – Efficient Long Video Generation
4:Encode $\mathcal{T},\mathcal{B}$ into conditioning vector $\mathcal{C}$
5:Initialize noise tensor $\mathbf{z}_{0}$
6:for $t=1$ to $T$ do
7:  $\mathbf{z}_{t}\leftarrow\text{FastDrive-DiT}(\mathbf{z}_{t-1},\mathcal{C})$ $\triangleright$ Apply diffusion steps with TeaCache
8:end for
9:Generated multi-view video frames $\{I_{i}^{1},\ldots,I_{i}^{T}\}_{i=1}^{N_{\text{views}}}\leftarrow\text{Decode}(\mathbf{z}_{T})$
10:Stage 2: FastRecon3D – Feed-Forward 3D Scene Reconstruction
11:for $t=1$ to $T$ do
12:  Collect frames $\{I_{i}^{t-\Delta},I_{i}^{t},I_{i}^{t+\Delta}\}_{i=1}^{N_{\text{views}}}$
13:  Estimate camera poses: $\{\mathbf{P}_{i}^{t}\}_{i=1}^{N_{\text{views}}}\leftarrow\text{PoseNet}(I_{i}^{t})$
14:  Estimate depth maps: $\{D_{i}^{t}\}_{i=1}^{N_{\text{views}}}\leftarrow\text{DepthNet}(I_{i}^{t})$
15:  Predict 3D Gaussian Parameters for each frame: $\mathcal{G}^{t}\leftarrow\text{GaussianNet}(I_{i}^{t},D_{i}^{t},\mathbf{P}_{i}^{t})$
16:end for
17:return Complete 3D dynamic scene $\mathcal{G}^{1:T}$
@@
DrivingForward [11] (640p) 21.67 0.727 0.259
DrivingForward [11] (228p) 21.76 0.767 0.194
Ours (w/ GT images) 23.71 0.733 0.285
Ours (w/ GEN images) 22.84 0.811 0.332
Table 1: Comparison of our method against prior feed-forward and optimization-based methods. The last two rows show novel view rendering performance with either GT or generated video input. All metrics are computed at frame $t$ given t1t{-}1 and t+1t{+}1 as inputs.
4 Experiments
4.1 Experimental Setup

### v2

@@ Abstract
We present DriveGen3D, a novel framework for generating high-quality and highly controllable dynamic 3D driving scenes that addresses critical limitations in existing methodologies. Current approaches to driving scene synthesis either suffer from prohibitive computational demands for extended temporal generation, focus exclusively on prolonged video synthesis without 3D representation, or restrict themselves to static single-scene reconstruction. Our work bridges this methodological gap by integrating accelerated long-term video generation with large-scale dynamic scene reconstruction through multimodal conditional control. DriveGen3D introduces a unified pipeline consisting of two specialized components: FastDrive-DiT, an efficient video diffusion transformer for high-resolution, temporally coherent video synthesis under text and Bird’s-Eye-View (BEV) layout guidance; and FastRecon3D, a feed-forward module that rapidly builds 3D Gaussian representations across time, ensuring spatial-temporal consistency. DriveGen3D enable the generation of long driving videos (up to $800\times 424$ at $12$ FPS) and corresponding 3D scenes, achieving state-of-the-art results while maintaining efficiency.
I Introduction
@@
The synthesis of 3D dynamic driving environments has emerged as a key research frontier in autonomous systems, driven by its wide-ranging applications in simulation, perception, and planning. While recent advances in video generation [1] and 3D scene reconstruction [2, 3] have made substantial progress, an integrated and efficient framework that unifies long-horizon video synthesis and large-scale 3D scene reconstruction under multi-modal control is still lacking.
@@
Existing methodologies typically address either temporal coherence in video generation or spatial fidelity in scene reconstruction, but not both—often requiring high computational cost or suffering from limited scalability. For example, state-of-the-art diffusion-based models like MagicDriveDiT [1] can produce high-resolution driving sequences but require up to 30 minutes to generate a single 20s video, making them impractical for real-time use.
@@
From the reconstruction perspective, optimization-based approaches [2, 4] are similarly time-consuming, often needing 30 minutes per scene. While recent feed-forward methods [5, 6, 3, 7, 8] have reduced reconstruction time to seconds, they remain limited in scale and rarely integrate with dynamic video generation pipelines.
@@
To bridge this gap, we propose DriveGen3D, an efficient and unified framework that integrates two specialized modules: FastDrive-DiT, an accelerated video diffusion transformer for high-resolution driving video generation, and FastRecon3D, a feed-forward reconstruction pipeline that builds dynamic 3D scenes from multi-view video frames in real-time.
To address the latency in video generation, we analyze the diffusion process and identify that applying caching strategies [9] exclusively to the conditional branch maximizes speedup without compromising layout control. Furthermore, we addressed the computational bottleneck of the attention layers. By profiling the attention maps, we apply precise 8-bit quantization [10, 11] specifically to the cross-view attention modules, reducing inference time by over $2\times$ while preserving visual fidelity.
FastRecon3D enables rapid driving scene reconstruction. Unlike prior methods that treat frames independently [5, 6] or optimize each scene individually [2], we propose a recursive, feed-forward and temporal-aware approach. By employing cross-attention to fuse features from past and future contexts, our model reconstructs 3D primitives that are both spatially accurate and temporally coherent. This effectively mitigates motion blur in dynamic elements and ensures consistent geometry across the synthesized sequence.
We conduct extensive experiments on the nuScenes dataset [12] demonstrating that DriveGen3D outperforms state-of-the-art baselines in both video quality and 3D reconstruction accuracy. Specifically, Our framework achieves a PSNR of 23.25, an SSIM of 0.820, and an LPIPS of 0.176 on generated inputs, significantly outperforming previous state-of-the-art methods. Moreover, DriveGen3D reduces the end-to-end generation time to under 6 minutes, including video synthesis and 3D scene reconstruction, marking a substantial improvement in efficiency. Our contributions can be summarized as follows:
@@
  • DriveGen3D achieves high-quality video generation and complete 3D reconstruction for large-scale street scenes within 6 minutes, significantly outperforming the state-of-the-art methods in terms of photorealism, scalability and efficiency.
II Related Work
II-A Video generation for driving scene
@@
Recent advancements in street view generation and autonomous driving scene synthesis have significantly improved the fidelity and controllability of synthetic data. MagicDrive [13] introduces a framework for street view generation with diverse 3D geometry controls, such as camera poses and 3D bounding boxes, enhancing tasks like BEV segmentation and 3D object detection through cross-view consistency. MagicDriveDiT [1] extends this by addressing high-resolution, long video generation for autonomous driving, leveraging flow matching and spatial-temporal conditional encoding to achieve superior scalability and control. DrivingDiffusion [14] synthesizes multi-view driving videos with precise layout control, ensuring cross-view and cross-frame consistency through a spatial-temporal diffusion framework. UniScene [15] unifies the generation of semantic occupancy, video, and LiDAR data, employing a progressive generation process to reduce complexity and improve downstream task performance.Together, these works advance the field of autonomous driving by providing scalable, controllable, and high-fidelity synthetic data generation frameworks.
II-B Reconstruction for driving scene
@@
The reconstruction of dynamic driving scenes [2] has emerged as a critical task in autonomous systems and immersive environment modeling. Contemporary approaches predominantly leverage Gaussian splatting-based representations due to their inherent balance between rendering efficiency and geometric expressiveness. Early methodologies in this domain adopted optimization-based paradigms, exemplified by works such as StreetGaussian [2] and DrivingGaussian [4]. These frameworks optimize scene representations per instance for specific street segments (typically under 100 meters in scale) through iterative refinement processes spanning approximately 30 minutes. Recent advancements have shifted toward feed-forward architectures [5, 6, 16, 17, 18, 19, 20] to enable rapid 3D reconstruction. Methods like PixelSplat [5] and MVSplat [6] employ large pretrained networks to directly infer Gaussian parameters from multi-view inputs, reducing reconstruction time from minutes to seconds. Though these approaches demonstrate generalizability across scenes, they often sacrifice reconstruction fidelity in geometrically complex regions or under sparse observational constraints. Parallel innovations address the temporal dimension of driving scenes: InfiniCube [21] extends the Scube [22] framework to 3D street generation, disentangling dynamic vehicles from static backgrounds via hybrid voxel-video control mechanisms. Drive3R [8] adapts the architecture of Spann3R [23] for per-frame 3D scene reconstruction through temporal consistency priors. DrivingRecon [7] mimics StreetGaussians [2] but replaces iterative optimization with feed-forward prediction. DrivingForward [3] enhances sparse-view reconstruction robustness by jointly learning pose estimation and depth prediction modules within its network architecture. These advancements collectively highlight two persistent limitations:
@@ III-A Overview
DriveGen3D is an integrated 3D driving scene generation framework composed of two key components as illustrated in Figure 1. The pipeline begins with FastDrive-DiT, which synthesizes high-resolution, temporally coherent driving videos under conditional guidance. These generated videos are then passed to FastRecon3D, which reconstructs dynamic 3D scenes in a feed-forward manner using temporal-aware Gaussian splatting. Together, these modules enable rapid and scalable 3D scene generation suitable for real-time simulation and autonomous driving applications.
The architecture of our efficient video diffusion module, dubbed FastDrive-DiT, integrates two targeted acceleration strategies to address the high inference cost of driving video generation, leading to a speedup of more than $2\times$. FastRecon3D proposes a temporally context-visible fusion module based on the static feed-forward framework, resulting in better temporally consistent per-frame Gaussians.
III-B FastDrive-DiT
@@
Generating 3D driving scenes, especially in the autonomous driving domain, is notoriously time-consuming due to the multi-view nature of the data. The video generation step is particularly costly because of the underlying diffusion process. For instance, MagicDriveDiT [1] can take up to 30 minutes to produce a video of driving scene. To address this inefficiency, we propose FastDrive-DiT, an enhanced and lightweight video diffusion model built on MagicDriveDiT with targeted acceleration strategies.
@@
Diffusion steps acceraleration. To accelerate the generation of long and high-resolution videos in MagicDriveDiT, we integrate TeaCache[9], a training-free caching approach that enhances the inference speed of diffusion models. TeaCache leverages timestep embeddings to modulate noisy inputs, approximating the fluctuating differences in model outputs across timesteps. By introducing a rescaling strategy, it refines these differences, enabling efficient output caching with minimal computational overhead. Experiments show that TeaCache achieves up to $4.41\times$ acceleration while maintaining visual quality, making it an effective solution for balancing speed and performance in video generation.
@@
In the original TeaCache[9] implementation, coefficients are computed from both the conditional and unconditional branches of the diffusion model. However, in our adaptation for MagicDriveDiT, we calculate coefficients exclusively from the conditional branch. This optimization not only reduces the forwarding time of TeaCache but also ensures no obvious degradation in performance, further enhancing the efficiency of the video generation process.
@@
Figure 2: Visualization of the multi-view reconstructed video from a generated 3D scene.
III-C FastRecon3D
While the aforementioned methods enable realistic driving scene generation, applications like simulation require complete 3D scene models. To enable rapid novel scene synthesis, we introduce FastRecon3D, a feed-forward reconstruction module that avoids costly optimization while preserving quality. Built upon temporal-aware Gaussian Splatting, FastRecon3D reconstructs the 3D geometry of each frame 3D geometry by leveraging both past and future contexts, enabling fast and consistent 3D scene generation.
@@ III-C FastRecon3D
$G_{i}^{t}=\{\mathcal{G}_{i}^{t}\}_{i=1}^{N_{t}}=\{(\mathbf{\mu}_{i},\mathbf{\Sigma}_{i},\mathbf{\alpha}_{i},\mathbf{c}_{i})\}_{i=1}^{N_{t}}.$ (1)
Recursively reconstruction from videos. We employ the decoder of DrivingForward [3] as the Gaussian Network to estimate Gaussian parameters, and use a pre-trained image encoder [25] to extract features.
After the frame generation of the $t+\Delta$ timestep, we extract all frames from time steps $t-\Delta$ to $t+\Delta$ to reconstruct the scene at time $t$. Our model first extracts features $\{F_{i}^{t}\}_{i=1}^{N_{t}}$ from each frame. For each timestep $t$, given multi-view features $\{F_{i}^{t}\}_{i=1}^{N_{t}}$ and temporal neighbors $\{F_{i}^{t\pm\Delta}\}_{i=1}^{N_{t}}$, we use cross-attention to fuse information from other frames into $\{F_{i}^{t}\}_{i=1}^{N_{t}}$:
$\{{F^{\prime}}_{i}^{t}\}_{i=1}^{N_{t}}=\text{Attention}(Q,K,V),$ (2)
$Q\leftarrow\{F_{i}^{t}\}_{i=1}^{N_{t}},\quad K,V\leftarrow\{F_{i}^{t\pm\Delta}\}_{i=1}^{N_{t}}.$ (3)
TABLE I: Comparison of our method against prior feed-forward and optimization-based methods. The last two rows show novel view rendering performance with either GT or generated video input.
@@
DrivingForward [3] 21.69 0.742 0.226
Ours (w/ GT images) 23.87 0.814 0.169
Ours (w/ GEN images) 23.25 0.820 0.176
Then the network predicts Gaussian parameters and save it as a 3D model:
$\mathcal{G}^{t}=\{\mu^{t},\Sigma^{t},\alpha^{t},c^{t}\}=\Psi\Bigl(\{{F^{\prime}}_{i}^{t-\Delta},\,{F^{\prime}}_{i}^{t},\,{F^{\prime}}_{i}^{t+\Delta}\}_{i=1}^{N_{t}}\Bigr).$ (4)
By leveraging both past and future context, this recursive reconstruction method effectively captures dynamic scene elements while maintaining high spatial fidelity. As a result, our approach produces complete 3D models in a matter of seconds, meeting the demanding requirements of simulation and other real-time applications without compromising on quality.
IV Experiments
IV-A Experimental Setup